# Towards an exact electronic quantum many-body treatment of Kondo correlation in magnetic impurities

### v1

@@ Results
As a result of the charge transfer, the local spin moments of the magnetic impurities differ from their isolated-atom values (Fig. 2b). For Mn, Fe, Co, and Ni, we predict $S=1.99,1.42,0.95,0.48$, about 0.5 lower than the isolated-atom spins. These $S$ values agree well with the available experimental data for Mn and Fe [Daybell1968, Haen1976, Joly2017]. Cr has the highest spin moment ($S=2.23$) among all $3d$ transition metal impurities, which deviates from the common understanding of this Kondo series [Nevidomskyy2009], but is also supported by experimental measurements that suggest Cr is a high-spin impurity ($S=5/2$[Haen1976, Sacramento1990]. From Eq. 3, the large moment of Cr in the Cu host suggests that the observed suppression of $T_{\mathrm{K}}$ relative to that of Fe arises from the same physical principles. However, it is also clear that the trends in $T_{\mathrm{K}}$ are not solely captured by the effective $S$ (see Supplementary Note 8 for orbital-resolved spin correlations).
Our predicted Kondo temperatures for all the studied elements are shown in Fig. 2c. We capture the exponential decay in the Kondo temperature moving from low-spin to high-spin impurities, in good agreement with the experimental estimates. The Kondo energy scales of Mn, Fe, Co, and Ni are found to be $e_{g}$-driven, while they are $t_{2g}$-driven for Cr, V, and Ti. With the exception of Cr and Mn, our predicted $T_{\mathrm{K}}$ is within a factor of 2$\sim$3 of at least one of the experimental estimates of $T_{\mathrm{K}}$. As our calculations do not include long-range Coulomb fluctuations (beyond those captured by the basis extending to neighbouring atoms), this suggests that they do not contribute significantly at this level of accuracy. On the other hand, Mn, which has the lowest $T_{\mathrm{K}}$, is a system where we could not achieve numerical convergence of our results to the parent basis. This is the likely origin of the larger quantitative discrepancy.
@@
Figure 3: Comparison between model and ab initio all-orbital simulations. (a) Orbital-resolved quasiparticle renormalization weights. (b) Predictions of Kondo temperatures against experimental values.
@@
Finally, to place the fidelity of our results in perspective, we check to see if the above predictions are easily captured within a standard application of the downfolded model approach. For this, we employ a five-orbital Anderson impurity model using the Kanamori Hamiltonian [Kanamori1963a, Georges2013] with screened Coulomb interaction and exchange parameters $U$ and $J$ taken from constrained random phase approximation (cRPA) calculations [Sasioglu2011]. (Note these parameters are also very close to previous parameters used in other model Kondo studies, e.g., for Co in Cu [Surer2012a, Valli2020]). We employed the same valence hybridization, bath discretization, and active-space DMRG solvers described above, and further details are given in Supplementary Note 9. Fig. 3 shows the estimated Kondo temperatures and quasiparticle renormalization obtained from this model. We find that the quasiparticle renormalization weights of most magnetic impurities are significantly overestimated compared to our all-orbital simulations. As a result, the predicted $T_{\mathrm{K}}$’s do not even show a clear exponential trend, with a large overestimation of $T_{\mathrm{K}}$. In addition, the relative $T_{\mathrm{K}}$ of Cr and Fe, one of the more interesting and subtle trends discussed above, is not reproduced in the model calculation. The quantitative improvement in the Kondo temperatures moving from the model calculations can be quite large: as much as an order of magnitude, or even more.
@@ Discussion.
In contrast to previous approaches, we achieved this accuracy by performing the quantum many-body simulations of the bare electronic problem with all orbitals, rather than within a downfolded model. This allows for a straightforward control of many aspects of converging our results towards an exact description of the phenomena. Within the zero-temperature electronic picture of our work, the physics of very long-range Coulomb fluctuations remain as an uncontrolled uncertainty, although these effects appear to be small on the scale of our results. However, our computational framework does not lie in opposition to model approaches. In particular, now that we can establish faithful physics within our framework, we can seek to extract traditional low-energy models from it. Moving forwards, the techniques used here to simulate Kondo physics are broadly applicable to more complicated correlated electron materials. The accuracy achieved here suggests that with continuing advances, we may move beyond qualitative descriptions of correlated electron phenomena, towards the quantitative simulation of observables directly measured by experiments.
@@ Acknowledgements
In the initial phase of this project performed at Caltech, T.Z. was supported by the Air Force Office of Scientific Research through the MURI program, Grant No. FA9550-18-1-0095. The development of integral infrastructure by Z.-H.C. was supported by the US Department of Energy, Office of Science, Basic Energy Sciences, through award DE-SC0018140. L.P.’s continuation of the work by T.Z. was supported by the Center for Molecular Magnetic Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Grant No. DE-SC0019330. T.Z. also acknowledges support from the Air Force Office of Scientific Research under award number FA9550-24-1-0096 for the work conducted at Yale. H.Z. (DMRG solver) was supported by the US National Science Foundation, through Grant No. CHE-2102505. Calculations were conducted in the Resnick High Performance Computing Center, supported by the Resnick Sustainability Institute at Caltech, as well as in the Yale Center for Research Computing.
@@
T.Z. and G.K.-L.C. designed the initial project and wrote the manuscript. T.Z., L.P., and H.Z. developed all-orbital DMFT method and code. T.Z., L.P., and Z.-H.C. performed ab initio calculations of Kondo systems. T.Z. carried out the model Hamiltonian calculations. H.Z. developed the code of DMRG impurity solver. G.K.-L.C. supervised the project. All authors contribute to the discussion of the results as well as the writing and editing of the manuscript.
@@
Methods, Figures S1-S6, Tables S1-S8.
@@
We incorporate the full impurity atom into the embedding problem. In the IAO+PAO basis, the $3d4s4p4d4f5s$ (def2-SVP) and $3d4s4p4d4f5s5p5d5f5g6s6p6d7s7p$ (cc-pVTZ) orbitals of the impurity atoms were treated in the many-body impurity solvers, while the $1s2s2p3s3p$ orbitals were frozen at the mean-field level and left out of the embedding problem. The fcDMFT code [Zhu2019, Zhu2020, Zhu2021c] was used to perform the all-orbital DMFT calculations. We employed the bare Coulomb interaction $(ij|kl)$ within the impurity as the two-particle interaction matrix in the embedding Hamiltonian in Eq. 1. The one-particle impurity interaction matrix in Eq. 1 is defined as
@@
where $\delta$ is the broadening factor taken to be $\delta=0.01$ Ha. Since Hamiltonian-based impurity solvers (that require a bath) were employed, we discretized the $3d4s$ block of the hybridization function on a non-uniform grid along the real frequency axis using a pole-merging direct discretization method [DeVega2015c] to obtain bath energies $\{\epsilon_{p}\}$ and impurity-bath couplings $\{V_{ip}\}$ in Eq. 1, and 49 bath orbitals were coupled to each of the $3d4s$ valence orbitals. Among these 49 bath orbitals, 5 orbitals were placed within a $\mu\pm 0.027$ eV energy window, 18 orbitals were placed within the $[\mu\pm 0.027,\mu\pm 0.6]$ eV energy window, and the remaining 26 orbitals were placed at a higher energy window of $[\mu\pm 0.6,\mu\pm 8.0]$ eV. In total, the embedding problem contains 22 (impurity) + 294 (bath) = 316 orbitals in the def2-SVP impurity basis calculations. The same bath discretization procedure was applied to the cc-pVTZ impurity basis calculation of the Fe impurity, resulting in an embedding problem of 59 (impurity) + 294 (bath) = 353 orbitals.
@@
where $C^{\mathrm{emb,act}}=C^{\mathrm{CISD}}C^{\mathrm{DMRG}}$. Here, $C^{\mathrm{CISD}}$ represents the rotation matrix from the full embedding space to the natural-orbital active space derived from CISD calculations, while $C^{\mathrm{DMRG}}$ represents the rotation matrix from the CISD active space to the natural-orbital active space derived from ground-state DMRG calculations. Finally, the local Green’s function of the impurity was calculated from the Dyson’s equation
@@
where $2D$ is the conduction electron bandwidth. For convenience, we set $D=1$. The flat-band hybridization on a logarithm grid was then discretized along the real axis to obtain $\{\epsilon_{k},V_{k}\}$:
@@
We first solved the SIAM Hamiltonian using the Hartree-Fock approximation, followed by a configuration interaction with single and double excitations (CISD) calculation. We then derived a natural-orbital-based active space by diagonalizing the CISD one-particle density matrix. We denote the $n$-electron $n$-orbital active space as ($n$e, $n$o). We then solved the active space using ground-state DMRG with bond dimension $M=1200$ and then dynamic DMRG with bond dimension $M=800$ [Zhai2021a, Zhai2023a]. To understand the accuracy of this active-space DMRG solver, we also solved the full 40-orbital SIAM problem using DMRG/DDMRG with the same bond dimensions.
Figure S1: Impurity density of states and self-energy in the single-impurity Anderson model calculated by full DMRG and active-space DMRG. $U=0.2$. (a) (4e, 4o) active space. (b) (8e, 8o) active space. (c) (12e, 12o) active space.
We first present the impurity density of states (DOS) and self-energy at $U=0.2$ in Fig. S1. It shows a sharp Kondo resonance peak and two broad Hubbard bands in the full DMRG calculation, which agrees quantitatively with previous NRG results [Bulla1998]. Compared to the full (40e, 40o) DMRG results (Fig. S1a), the (4e, 4o) active-space result shows multiple spurious oscillations in the DOS and self-energy. However, we point out that even with a very small (4e, 4o) active space, the impurity solver correctly predicts the shape of the Kondo resonance around the Fermi level, which is also confirmed in the self-energy comparison around $\omega=0$. We find that the accuracy improves rapidly when the active space is increased from (4e, 4o) to (8e, 8o) and (12e, 12o). At the (12e, 12o) level, the DOS and self-energy from active-space DMRG are almost indistinguishable from the full DMRG results.
Figure S2: Quasiparticle renormalization weight $Z$ for increasing sizes of active space in $U=0.1,0.2,0.3$ SIAM models. The full (40e, 40o) DMRG calculated $Z$ are shown as dashed lines.
@@
$Z=\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}.$ (S9)
We find that at moderate interaction strength ($U=0.1$), the active-space DMRG is already accurate at the (12e, 12o) level, predicting $Z=0.22$ vs. the full DMRG result $Z=0.20$. At stronger interaction strength ($U=0.2$), the (12e, 12o) active-space DMRG predicts $Z=0.035$, an overestimation compared to the full DMRG result $Z=0.018$. In summary, for stronger interactions ($U=0.2$ and $U=0.3$), a larger active space than (12e, 12o) is needed to achieve a completely quantitative description of the quasiparticle renormalization. Nevertheless, at all active space sizes, the active-space DMRG solver correctly predicts the exponential decay in $Z$ as $U$ increases from 0.1 to 0.3. We point out that it is difficult to directly infer the accuracy of the active-space DMRG in the ab initio Kondo simulations from the SIAM calculations alone because of the existence of multiple $3d$ orbitals and the inclusion of all other impurity valence and virtual orbitals into the embedding problem. However, the convergence of the quantities in the SIAM calculation with the different components of the calculation procedure (which are the same in the SIAM and ab initio calculations) supports the convergence checks that we discuss in the main text.

### v3

@@ Results
As a result of the charge transfer, the local spin moments of the magnetic impurities differ from their isolated-atom values (Fig. 2b). For Mn, Fe, Co, and Ni, we predict $S=1.99,1.42,0.95,0.48$, about 0.5 lower than the isolated-atom spins. These $S$ values agree well with the available experimental data for Mn and Fe [Daybell1968, Haen1976, Joly2017]. Cr has the highest spin moment ($S=2.23$) among all $3d$ transition metal impurities, which deviates from the common understanding of this Kondo series [Nevidomskyy2009], but is also supported by experimental measurements that suggest Cr is a high-spin impurity ($S=5/2$[Haen1976, Sacramento1990].
Taking into account the quantitative differences in the observed $S$ in the Kondo systems and the isolated atoms, our results partially confirm a Hund’s driven suppression mechanism. With the exception of the relative ordering of $T_{\mathrm{K}}(\mathrm{Cr})$ and $T_{\mathrm{K}}(\mathrm{Mn})$, higher observed $S$ indeed correlates with lower $T_{\mathrm{K}}$. For example, accounting for the charge transfer to Fe and Cr that modifies their effective spin, we can predict the observed relation $T_{\mathrm{K}}(\mathrm{Cr})<T_{\mathrm{K}}(\mathrm{Fe})$. The Hund’s picture can further be confirmed by examining the nature of the lowest triplet excited states which are associated with the destruction of the Kondo singlet. As seen in Fig. 2d, for most of the elements, the main character of the excitation is a coherent spin-flip of all the (close to) singly occupied $d$ orbitals, as expected in the limit of strong Hund’s coupling.
The exception to this is Mn. Note that from the computed effective spin, the Hund’s mechanism would predict Cr to have a lower $T_{\mathrm{K}}$ than Mn, contrary to the experimental findings. However, after charge transfer, the Mn ion has an excess (0.66) electron in the $t_{2g}$ shell, which experiences an unfavourable repulsion. We find that the lowest triplet excitation corresponds to a partial transfer of an electron from $t_{2g}$ to the bath, accompanied by a spin-flip (see Supplementary Note 13). The low energy of this triplet state recalls the double-exchange mechanism [anderson1955considerations], which favours same-spin alignment to enable charge delocalization. In fact, in our calculations, other than for Cr (which has a close to exactly half-filled $d$-shell when embedded in Cu) this type of excitation occurs at low energy in all the elements, but is especially low in energy for Mn. The inclusion of orbital degrees of freedom is thus an important part of the systematic trends in the Kondo temperatures, not only through the effective spin, but also through fluctuations.
Some additional insight is provided by examining the detailed contributions to Hewson’s formula. Because it starts from a Fermi liquid it naturally captures some effects omitted in the Kondo model, such as the effect of charge fluctuations and the crystal field splitting. Generally, the lower of the $T_{\mathrm{K}}(t_{2g})$ and $T_{\mathrm{K}}(e_{g})$ estimates corresponds to the shell with the most unpaired electron character, consistent with a Kondo model where the impurity is replaced by a spin. However, for impurities where both $t_{2g}$ and $e_{g}$ have significant unpaired electron character (such as in Cr) the Hewson formula further identifies which shell gives rise to the lower Kondo temperature. We find that the Kondo energy scales of Mn, Fe, Co, and Ni are associated with $e_{g}$, while they are associated with $t_{2g}$ for Cr, V, and Ti. As $\Delta(0)$ only varies slightly across the series (from $-0.59$ eV (Ti) to $-0.43$ eV (Fe) for the $t_{2g}$ orbital and from $-0.64$ eV (Ti) to $-0.38$ eV (Ni) for the $e_{g}$ orbital), the trend is generated primarily by the QP renormalization weight, which contains both the Hund’s coupling mechanism as well as the effect of charge fluctuations.
@@
Figure 3: Comparison of ab initio all-orbital simulations with model simulations using either the same chemical potentials (“Model”) or fitted chemical potentials to match the total $3d$ orbital occupancy of the all-orbital simulations (“Model (fit $\mu$)”). (a) Orbital-resolved quasiparticle renormalization weights. (b) Predicted Kondo temperatures compared to the experimental estimates.
@@
Finally, to place the fidelity of our results in perspective, we check to see if the above predictions are easily captured within a standard application of the downfolded model approach. For this, we employ a five-orbital Anderson impurity model using the Kanamori Hamiltonian [Kanamori1963a, Georges2013] with screened Coulomb interaction and exchange parameters $U$ and $J$ taken from constrained random phase approximation (cRPA) calculations [Sasioglu2011]. (Note these parameters are also very close to previous parameters used in other model Kondo studies, e.g., for Co in Cu [Surer2012a, Valli2020]). We employed the same valence hybridization, bath discretization, and active-space DMRG solvers described above. The model was tested using both the same chemical potentials as in the all-orbital calculations, and refitted ones to match the total $3d$ orbital occupancies in the all-orbital results. Further details and analysis are given in Supplementary Notes 10 and 11. Fig. 3 shows the estimated Kondo temperatures and quasiparticle renormalizations obtained from this model. We find that the quasiparticle renormalization weights of most magnetic impurities are significantly overestimated compared to our all-orbital simulations. As a result, the predicted $T_{\mathrm{K}}$’s do not even show a clear exponential trend, with a large overestimation of $T_{\mathrm{K}}$. In addition, the relative $T_{\mathrm{K}}$’s of Cr, Mn, and Fe, are not reproduced using either the same chemical potential as in the ab initio all-orbital calculations, or the refitted chemical potential. The quantitative improvement in the Kondo temperatures moving from the model to ab initio calculations is as much as 2 orders of magnitude. Our standard model includes the density-density, spin-flip, and pair-hopping interactions in rotationally invariant form. In principle, a model that allows for all possible terms in the $3d$ shell, together with their full frequency dependence, should yield better estimates of the Kondo temperatures [Valli2020]. However, determining such a large number of downfolded parameters is challenging.
@@ Discussion.
In contrast to previous approaches, we achieved this accuracy by performing the quantum many-body simulations of the bare electronic problem with all orbitals, rather than within a downfolded model. This allows for a straightforward control of many aspects of converging our results towards an exact description of the phenomena. Within the zero-temperature electronic picture of our work, the physics of very long-range Coulomb fluctuations remain as an uncontrolled uncertainty, although these effects appear to be small on the scale of our results. However, our computational framework does not lie in opposition to model approaches. In particular, we showed how, after establishing the faithfulness of the physics within our framework, we can extract understanding in terms of traditional low-energy pictures. Moving forwards, the accuracy achieved here suggests that with continuing advances, we may move beyond qualitative descriptions of correlated electron phenomena, towards a systematically improvable simulation of observables directly measured by experiments.
@@ Acknowledgements
In the initial phase of this project performed at Caltech, T.Z. was supported by the Air Force Office of Scientific Research through the MURI program, Grant No. FA9550-18-1-0095. The development of integral infrastructure by Z.-H.C. was supported by the US Department of Energy, Office of Science, Basic Energy Sciences, through award DE-SC0018140. L.P.’s continuation of the work by T.Z. was supported by the Center for Molecular Magnetic Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Grant No. DE-SC0019330. T.Z. acknowledges support from the Air Force Office of Scientific Research under award number FA9550-24-1-0096 for the work conducted at Yale. H.Z. (DMRG solver) was supported by the US National Science Foundation, through Grant No. CHE-2102505. GKC is a Simons Investigator in Physics. Calculations were conducted in the Resnick High Performance Computing Center, supported by the Resnick Sustainability Institute at Caltech, as well as in the Yale Center for Research Computing.
@@
T.Z., L.P., and G.K.-L.C. designed the initial project and wrote the manuscript. T.Z., L.P., and H.Z. developed the all-orbital DMFT method and code. T.Z., L.P., and Z.-H.C. performed ab initio calculations of Kondo systems. T.Z. carried out the model Hamiltonian calculations. H.Z. developed the code of the DMRG impurity solver. R.C. provided additional theoretical analysis. G.K.-L.C. supervised the project. All authors contribute to the discussion of the results as well as the writing and editing of the manuscript.
@@
Methods, Figures S1-S13, Tables S1-S13.
@@
We incorporate the full impurity atom into the embedding problem. In the IAO+PAO basis, the $3d4s4p4d4f5s$ (def2-SVP) and $3d4s4p4d4f5s5p5d5f5g6s6p6d7s7p$ (cc-pVTZ) orbitals of the impurity atoms were treated in the many-body impurity solvers, while the $1s2s2p3s3p$ orbitals were frozen at the mean-field level and left out of the embedding problem. The fcDMFT code [Zhu2019, Zhu2020, Zhu2021c] was used to perform the all-orbital calculations. We employed the bare Coulomb interaction $(ij|kl)$ within the impurity as the two-particle interaction matrix in the embedding Hamiltonian in Eq. 1. The one-particle impurity interaction matrix in Eq. 1 is defined as
@@
where $\delta$ is the broadening factor taken to be $\delta=0.01$ Ha. Since Hamiltonian-based impurity solvers (that require a bath) were employed, we discretized the $3d4s$ block of the hybridization function on a non-uniform grid along the real frequency axis using a pole-merging direct discretization method [DeVega2015c] to obtain bath energies $\{\epsilon_{p}\}$ and impurity-bath couplings $\{V_{ip}\}$ in Eq. 1, and 49 bath orbitals were coupled to each of the $3d4s$ valence orbitals. Among these 49 bath orbitals, 5 orbitals were placed within a $\mu\pm 0.027$ eV energy window, 18 orbitals were placed within the $[\mu\pm 0.027,\mu\pm 0.6]$ eV energy window, and the remaining 26 orbitals were placed at a higher energy window of $[\mu\pm 0.6,\mu\pm 8.0]$ eV. In total, the embedding problem contained 22 (impurity) + 294 (bath) = 316 orbitals in the def2-SVP impurity basis calculations. The same bath discretization procedure was applied to the cc-pVTZ impurity basis calculation of the Fe impurity, resulting in an embedding problem of 59 (impurity) + 294 (bath) = 353 orbitals.
@@
where $C^{\mathrm{emb,act}}=C^{\mathrm{CISD}}C^{\mathrm{DMRG}}$. Here, $C^{\mathrm{CISD}}$ represents the rotation matrix from the full embedding space to the natural-orbital active space derived from CISD calculations, while $C^{\mathrm{DMRG}}$ represents the rotation matrix from the CISD active space to the natural-orbital active space derived from ground-state DMRG calculations. Finally, the local Green’s function of the impurity was calculated from Dyson’s equation
@@
where $2D$ is the conduction electron bandwidth. For convenience, we set $D=1$. The flat-band hybridization on a logarithmic grid was then discretized along the real axis to obtain $\{\epsilon_{k},V_{k}\}$:
@@
We first solved the SIAM Hamiltonian using the Hartree-Fock approximation to determine an initial set of orbitals. Within this (untruncated basis) we subsequently carried out a ground-state configuration interaction with single and double excitations (CISD) calculation or a ground-state DMRG calculation. Using these, we then derived CISD and DMRG natural-orbital-based active spaces by diagonalizing the one-particle density matrices. We denote the $n$-electron $n$-orbital active space as ($n$e, $n$o). We then solved within the active space using ground-state DMRG with bond dimension $M=1200$ and then dynamic DMRG with bond dimension $M=800$ [Zhai2021a, Zhai2023a]. To understand the accuracy of this active-space DMRG solver, we also solved the full 40-orbital SIAM problem using DMRG/DDMRG with the same bond dimensions.
Figure S1: Impurity density of states and self-energy in the single-impurity Anderson model calculated by full DMRG and active-space DMRG (in CISD natural orbital basis). $U=0.2$. (a) (4e, 4o) active space. (b) (8e, 8o) active space. (c) (12e, 12o) active space.
We first present the impurity density of states (DOS) and self-energy at $U=0.2$ in Fig. S1. It shows a sharp Kondo resonance peak and two broad Hubbard bands in the full DMRG calculation, which agrees quantitatively with previous NRG results [Bulla1998]. Compared to the full (40e, 40o) DMRG results (Fig. S1a), the CISD (4e, 4o) active-space result shows multiple spurious oscillations in the DOS and self-energy. However, we point out that even with this very small CISD (4e, 4o) active space, the impurity solver correctly predicts the shape of the Kondo resonance around the Fermi level, which is also confirmed in the self-energy comparison around $\omega=0$. We find that the accuracy improves rapidly when the CISD active space is increased from (4e, 4o) to (8e, 8o) and (12e, 12o). At the (12e, 12o) level, the DOS and self-energy from active-space DMRG are almost indistinguishable from the full DMRG results.
Figure S2: Quasiparticle renormalization weight $Z$ for increasing sizes of active space in $U=0.1,0.2,0.3$ SIAM models. The active spaces consist of CISD natural orbitals (“CISD-NO”) or DMRG natural orbitals (“DMRG-NO”). The full (40e, 40o) DMRG calculated $Z$ and analytic (exact) $Z$ values are shown as horizontal dashed lines.
@@
$Z=\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}$ (S9)
using the CISD and DMRG active spaces, denoted “CISD-NO” and “DMRG-NO,” respectively. At moderate interaction strength ($U=0.1$), the active-space DMRG predicts $Z=0.21$ at the CISD (16e, 16o) level, in excellent agreement with full DMRG $Z=0.20$ and the exact $Z=0.21$ [Hewson1993a]. The exact $Z$ is analytically calculated from Hewson’s formula $Z=4T_{\mathrm{K}}/\pi\Delta_{0}$ (with $\Delta_{0}=0.015$) and the $T_{\mathrm{K}}$ for large $\frac{U}{\Delta_{0}}$ [Hewson1993a]
$T_{\mathrm{K}}=\sqrt{\frac{U\Delta_{0}}{2}}\exp\left(-\frac{\pi U}{8\Delta_{0}}+\frac{\pi\Delta_{0}}{2U}\right).$ (S10)
However, at stronger interaction strengths, the CISD active spaces become less accurate and overestimate $Z$. In contrast, the DMRG active spaces are more accurate, and even with the small (8e, 8o) active space, the computed $Z$ values are reasonably close to the full DMRG and exact results: at $U=0.2$ ($U=0.3$), the active-space DMRG predicts $Z=0.012$ ($Z=0.0017$), while the full DMRG predicts $Z=0.019$ ($Z=0.0019$) and the exact result is $Z=0.020$ ($Z=0.0017$). With the DMRG (16e, 16o) active space, the computed $Z$ values are almost indistinguishable from the exact results on the scale of Fig. S2. This confirms the effectiveness of an active-space strategy to reduce the number of orbitals needed to represent the properties of the impurity.
Table S1: Orbital occupancy cut-off thresholds for DMRG occupied and virtual natural orbitals to obtain active spaces of 6-16 orbitals in $U=0.2$ SIAM model.
Active space size Threshold (occupied) Threshold (virtual)
6o $2\times 10^{-3}$ $2\times 10^{-3}$
8o $4\times 10^{-4}$ $4\times 10^{-4}$
10o $9\times 10^{-5}$ $8\times 10^{-5}$
16o $2\times 10^{-6}$ $8\times 10^{-7}$
Table S2: Orbital occupancy cut-off thresholds for occupied and virtual natural orbitals to obtain the largest active spaces used in dynamical and finite temperature calculations (around 28 orbitals) in ab initio calculations of various impurity elements.
Element Threshold (occupied) Threshold (virtual)
Ti $4\times 10^{-5}$ $4\times 10^{-4}$
V $2\times 10^{-4}$ $7\times 10^{-4}$
Cr $4\times 10^{-4}$ $3\times 10^{-3}$
Mn $2\times 10^{-4}$ $4\times 10^{-3}$
Fe $1\times 10^{-4}$ $4\times 10^{-3}$
Co $1\times 10^{-4}$ $1\times 10^{-3}$
Ni $2\times 10^{-5}$ $2\times 10^{-3}$
We cannot rigorously infer the accuracy of the active-space DMRG in the ab initio Kondo simulations from the SIAM calculations because of the different Hamiltonian. However, we can assign a rough correspondence to the SIAM problem by dividing the number of active space orbitals used in the ab initio calculation by the number of effective singly occupied orbitals in the impurity atom, and using $Z$ in the ab initio calculations to define a corresponding effective $U$. The most difficult systems are those with the lowest Kondo temperatures, namely Cr, Mn, and Fe, which have between 3-5 singly occupied orbitals. For Cr and Fe, the largest active space used in the computation of $Z$ had 28 orbitals, thus we infer an effective SIAM active space size of 6-10 orbitals. Similar correspondence can be obtained by comparing the natural orbital cut-off thresholds used to obtain active spaces in the ab initio calculations and in the SIAM model (Tables S1 and S2). The cut-off thresholds in ab initio calculations are on the order of $10^{-4}$ or $10^{-3}$, which are in the range of the thresholds of SIAM active spaces of 6-10 orbitals. The renormalization factors of Cr and Fe computed in the largest active space were 0.023 and 0.026, respectively. These $Z$ values correspond roughly to $U=0.2$, where with this effective (CISD) active space size in the SIAM problem we expect $Z$ to be overestimated by $\sim 3$, which is similar to the observed overestimation of $T_{\mathrm{K}}$ with respect to the experimental estimates in the ab initio calculations. For Mn, where we could only use a 22-orbital active space (see below) and $Z\approx 0.006$, corresponding to $U>0.2\sim 0.3$, we expect $Z$ to be overestimated by an order of magnitude. This also roughly agrees with the larger overestimation of $T_{\mathrm{K}}$ for Mn. The convergence of the SIAM calculations can thus be seen to be consistent with our ab initio results, and further provide an estimate of the required resources for more accurate ab initio calculations in the future.